# Open-MAGVIT2: An Open-Source Project Toward Democratizing Auto-regressive Visual Generation

### v1

@@ 1 Introduction
In this work, we push forward the auto-regressive visual generation in two folds: 1) Replication of the visual tokenizer: We re-implement the advanced Lookup-Free Quantizer proposed by MAGVIT-v2. To our best knowledge, our open-source replication achieves the closest reconstruction performance stated in MAGVIT-v2 (1.18 vs. 1.15 rFID on ImageNet 128$\times$128) and outperforms all other methods on the hallmark Imagenet benchmark (Deng et al., 2009). 2) Integrating a super-large codebook with AR visual generation: Instead of simply following MAGVIT-v2 that leverages the vision-oriented design (i.e., mask generative methods (Chang et al., 2022) for visual synthesis), we seek to exploit the potential of such a large codebook in vanilla auto-regressive generation. To assist auto-regressive models in predicting with a super-large vocabulary, we factorize it into two sub-vocabulary of different sizes by asymmetric token factorization, and further introduce “next sub-token prediction” to enhance sub-token interaction for better generation quality. Our experiments on the standard visual generation dataset ImageNet suggest that, with the powerful tokenizer, the plain auto-regressive model exhibits superiority and scalability.
@@ 3.1 Dataset and Metrics
The training of the visual tokenizer and auto-regressive transformer are both on ImageNet (Deng et al., 2009). Specifically, we train the tokenizer in $128\times 128$ and $256\times 256$ resolutions.
For visual reconstruction, the reconstruction-FID, denoted as rFID (Heusel et al., 2017), codebook utilization, the use percentage of codes, and PSNR on ImageNet 50k validation set are adopted to measure the quality of reconstructed images. Simultaneously, we measure the quality of image generation by the prevalent metrics FID, IS (Salimans et al., 2016) and Precision/Recall (Kynkäänniemi et al., 2019).
@@ 3.2 Implementations Details
Open-MAGVIT2 follows the same architecture of the visual tokenizer proposed in (Yu et al., 2024a). For computational efficiency, we remove the gradient penalty loss, and adopt PatchGAN (Isola et al., 2017) as the discriminator instead of StyleGAN (Karras et al., 2019). All models corresponding to different resolutions are trained with similar settings: an initial $41e-$ 4 learning rate, an Adam Optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $256$ batch size from $270$ to $350$ epochs, a combination of reconstruction, GAN, perceptual (Zhang et al., 2018), entropy penalty (Yu et al., 2024a), commitment losses, LeCAM regularization (Tseng et al., 2021) for training stability, and $32$$\times$ Nvidia V100 / Ascend 910B with Pytorch.
@@
As illustrated, we propose asymmetric token factorization to assist the auto-regressive transformer models in making the precise prediction with a large codebook. Note that, we empirically set $M=2$ and $k_{1}=6$, $k_{2}=12$. Since our main focus is on democratizing scalable auto-regressive visual generation, the plain auto-regressive transformer is utilized while the techniques that introduce inductive bias such as AdaLn (Karras et al., 2020) are excluded. Specifically, we adopt the Llama-based (Touvron et al., 2023) architecture (e.g., RoPE (Su et al., 2024), SwiGLU (Shazeer, 2020), RMSNorm (Zhang et al., 2022) technique, each of which has been proven effective in (Sun et al., 2024)). The class embedding which is indexed from a set of learnable embeddings serves as the start token. Open-MAGVIT2 follows the simple scaling principle proposed in (Sun et al., 2024), which is in Tab. 1. All models are trained with similar settings: a base learning rate of $41e-$ 4 per $256$ batch size, an AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = $25e-$ 2, a total $768$ batch size and 300 $\sim$ 350 training epochs, gradient clipping of $1.0$, $0.1$ dropout rate for input embedding, FFN module and conditional embedding, $32$ $\sim$ 96 $\times$ Nvidia V100 / Ascend 910B for different scales of the model with Pytorch.
@@ 3.3 Main Results
Open-MAGVIT2 2D $16\times 16$ 16 $256\times 256$ 262144 1.17 21.90 100%
@@
Open-MAGVIT2 2D $32\times 32$ 8 $128\times 128$ 262144 0.34 26.19 100%

### v2

@@ 1 Introduction
In this work, we push forward the auto-regressive visual generation in two folds: 1) Replication of the visual tokenizer: We re-implement the advanced Lookup-Free Quantizer proposed by MAGVIT-v2. To our best knowledge, our open-source replication achieves the closest reconstruction performance stated in MAGVIT-v2 (1.18 vs. 1.15 rFID on ImageNet 128$\times$128) and outperforms all other methods on the hallmark Imagenet benchmark (Deng et al., 2009). Furthermore, we provide the visual tokenizer for text-conditional image generation with improved representational capacity by pretraining it on large-scale general-domain datasets. 2) Integrating a super-large codebook with AR visual generation: Instead of simply following MAGVIT-v2 that leverages the vision-oriented design (i.e., mask generative methods (Chang et al., 2022) for visual synthesis), we seek to exploit the potential of such a large codebook in vanilla auto-regressive generation. To assist auto-regressive models in predicting with a super-large vocabulary, we factorize it into two sub-vocabulary of different sizes by asymmetric token factorization, and further introduce “next sub-token prediction” to enhance sub-token interaction for better generation quality. Our experiments on the standard visual generation dataset ImageNet suggest that, with the powerful tokenizer, the plain auto-regressive model exhibits superiority and scalability.
@@ 2.2 Visual Tokenizer
Figure 3: Reconstruction comparison with different tokenizers . We compare LlamaGen (Sun et al., 2024), Showo (Xie et al., 2024), Cosmos (Agarwal et al., 2025) and our models. The bottom of each case illustrates that the tokenizer pretrained on large-scale datasets gains more superiority in facial and textual scenarios than ones trained on imagenet. Moreover, our pretrained visual tokenizer achieves better reconstruction soundness compared with LlamaGen, Show-o and Cosmos. (Best viewed with zooming in.)
@@ 3.1 Dataset and Metrics
Class-Conditional Image Generation.
The training of the visual tokenizer and auto-regressive transformer are both on ImageNet (Deng et al., 2009). Specifically, we train the tokenizer in $128\times 128$ and $256\times 256$ resolutions. For visual reconstruction, the reconstruction-FID, denoted as rFID (Heusel et al., 2017), codebook utilization, the use percentage of codes, and PSNR on ImageNet 50k validation set are adopted to measure the quality of reconstructed images. Simultaneously, we measure the quality of image generation by the prevalent metrics FID, IS (Salimans et al., 2016) and Precision/Recall (Kynkäänniemi et al., 2019).
Text-Conditional Image Generation.
Aiming to serve text-conditional image generation, we pretrain the tokenizer on large-scale general-domain datasets, i.e., 1) General: LAION-COCO (LAION, 2022), CC12M (Changpinyo et al., 2021) and CC3M (Sharma et al., 2018). 2) High-quality: LAION-aesthetics-12M [^1] https://huggingface.co/datasets/dclure/laion-aesthetics-12m-umap, LAION-aesthetics (Schuhmann & Beaumont, 2022), JourneyDB (Pan et al., 2023) and LAION-HD [^2] https://huggingface.co/datasets/yuvalkirstain/laion-hd-subset. Following (Agarwal et al., 2025), we evaluate our visual tokenizer by the zero-shot performance on MS-COCO (Lin et al., 2014) and Imagenet with rFID, PSNR, SSIM.
@@ 3.2 Implementations Details
Open-MAGVIT2 follows the same architecture of the visual tokenizer proposed in (Yu et al., 2024a). The visual tokenizer for different purposes, e.g., class-conditional and text-conditional image generation, is trained with similar settings. Specifically, for computational efficiency, we remove the gradient penalty loss, and adopt PatchGAN (Isola et al., 2017) as the discriminator instead of StyleGAN (Karras et al., 2019). All models corresponding to different resolutions are trained with similar settings: an initial $41e-$ 4 learning rate, an Adam Optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $256$ batch size from $270$ to $350$ epochs (1500k steps for text-conditional image generation), a combination of reconstruction, GAN, perceptual (Zhang et al., 2018), entropy penalty (Yu et al., 2024a), commitment losses, LeCAM regularization (Tseng et al., 2021) for training stability, and $32$$\times$ GPU / NPU with Pytorch.
@@
As illustrated, we propose asymmetric token factorization to assist the auto-regressive transformer models in making the precise prediction with a large codebook. Note that, we empirically set $M=2$ and $k_{1}=6$, $k_{2}=12$. Since our main focus is on democratizing scalable auto-regressive visual generation, the plain auto-regressive transformer is utilized while the techniques that introduce inductive bias such as AdaLn (Karras et al., 2020) are excluded. Specifically, we adopt the Llama-based (Touvron et al., 2023) architecture (e.g., RoPE (Su et al., 2024), SwiGLU (Shazeer, 2020), RMSNorm (Zhang et al., 2022) technique, each of which has been proven effective in (Sun et al., 2024)). The class embedding which is indexed from a set of learnable embeddings serves as the start token. Open-MAGVIT2 follows the simple scaling principle proposed in (Sun et al., 2024), which is in Tab. 1. All models are trained with similar settings: a base learning rate of $41e-$ 4 per $256$ batch size, an AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay = $25e-$ 2, a total $768$ batch size and 300 $\sim$ 350 training epochs, gradient clipping of $1.0$, $0.1$ dropout rate for input embedding, FFN module and conditional embedding, $32$ $\sim$ 96 $\times$ GPU / NPU for different scales of the model with Pytorch.
@@ 3.3 Main Results
Open-MAGVIT2 2D $16\times 16$ 16 $256\times 256$ 262144 1.17 22.64 100%
@@
Open-MAGVIT2 2D $32\times 32$ 8 $128\times 128$ 262144 0.34 27.02 100%
@@
Visual Reconstruction.
Furthermore, to facilitate the development of autoregressive text-conditional image generation, we provide an improved version of our tokenizers by pretraining on large-scale image-text datasets. As illustrated in Fig. 3, our tokenizers achieve state-of-the-art performance compared to concurrent methods such as Cosmos (Agarwal et al., 2025) in terms of zero-shot reconstruction on both ImageNet and COCO datasets. It is worth noting that some recent efforts in residual tokenization (Qu et al., 2024; Han et al., 2024) can achieve better results, but are not listed here because residual techniques are orthogonal and compatible with quantization methods such as VQ, LFQ, and FSQ.
@@
Table 5: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017. The tokenizers are trained with large-scale general-domain datasets and aim to serve text-conditional image generation. The results are reported under the same setup. $\dagger$ indicates that LlamaGen loads the model initially trained on Imagenet while the others are training from scratch, i.e., MS-COCO and Imagenet-1k are excluded from training data. Text in gray signifies the metrics adopted from Cosmos (Agarwal et al., 2025) report.
Method Quantizer Training Ratio Codebook MS-COCO 2017 Imagenet-1k
[…]
Resize $256\times 256$
LlamaGen (Sun et al., 2024) VQ 70M 16 16384 8.40 20.28 0.55 2.47 20.65 0.54
Show-o (Xie et al., 2024) LFQ 35M 16 8192 9.26 20.90 0.59 3.50 21.34 0.59
Cosmos (Agarwal et al., 2025) FSQ - 16 64000 11.97 19.22 0.48 4.57 19.93 0.49
Open-MAGVIT2 LFQ 100M 16 16384 7.93 22.21 0.62 2.55 22.21 0.62
Open-MAGVIT2 LFQ 100M 16 262144 6.76 22.31 0.65 1.67 22.70 0.64
Original Resolution
Cosmos (Agarwal et al., 2025) FSQ - 16 64000 7.51 20.45 0.52 1.93 20.56 0.51
Cosmos (Agarwal et al., 2025) FSQ - 16 64000 7.23 20.45 0.53 2.52 20.49 0.52
Open-MAGVIT2 LFQ 100M 16 16384 6.65 21.61 0.57 1.39 21.74 0.56
Open-MAGVIT2 LFQ 100M 16 262144 5.10 22.18 0.60 0.78 22.24 0.59